Find the minimum sample size n needed to estimate mu for the given values of​ c, sigma​, and E. cequals0.90​, sigmaequals7.5​, and Eequals2 Assume that a preliminary sample has at least 30 members.
Answer #1

Solution :

Given that,

standard deviation = \sigma = 7.5

margin of error = E = 2

At 90% confidence level the z is ,

\alpha = 1 - 90% = 1 - 0.90 = 0.10

\alpha / 2 = 0.10 / 2 = 0.05

Z\alpha/2 = Z0.05 = 1.645

Sample size = n = ((Z\alpha/2 * \sigma ) / E)2

= ((1.645 * 7.5 ) / 2)2

= 37.99

= 38

Sample size = 38